invest in a stock portollo. Your chaices ane Stock X with an expected retum of 16 pecent and Stock Y with an expected retum of 10 percant Assume your gol is te oae aporle with aedum at 1aso per How much money wil you invest in Stook X and Stock Y? (De nat round intermediats cakulations and round your answers to 2 decimal places, .g 32.18)
Answer #1

assume invest in stock X=w

then invest in stock Y=1-w

expected return=13.50%

w*16%+(1-w)*10%=13.50%

w*6%+10%=13.50%

w=3.50%/6%=0.5833

invest in stock X=16000*0.5833=9332.80

invest in stock Y=16000*(1-0.5833)=6667.20

the above is answer..